1. Market research shows that for an increase in price by $1/kg, quantity demand decreased by 250 Kg.
  2. In addition, when the price is $1/kg, 1500 Kg of apple is demanded in the country

what is Qdx?


1
Expert's answer
2021-12-29T12:31:57-0500

"Edp=\frac{\Delta Q}{\Delta P}\times\frac{P}{Q}=\frac{250}{1}\times\frac{1}{1500}=0.17"

The elasticity of demand relative to price is the limit of the ratio of the relative increment in demand to the relative increment in price, provided that the increment in price tends to zero.